OpenStudy (anonymous):

Solve by using the perfect squares method. x power of 2 + 8x + 16 = 0 2 -2 4 -4

OpenStudy (lalaly):

(x+4)^2 =0 x=-4 or u can use the quadratic formula

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

x+4=0 x+4=0 -4 -4 -4 -4 x=-4 x=-4

OpenStudy (anonymous):